# Optimal Algorithm for Multiple-Processor Multitask Scheduling

### v1

@@ 6 Experimental Results
7 Conclusion
In this work, we study the multitask scheduling problem, for which the optimal algorithms and tight analyses remain widely open for almost all settings. We propose $\mathsf{NP}$-$\mathsf{SRPT}$ algorithm, which achieves a competitive ratio that is order optimal when the number of machines is constant. Another appealing and more important property of $\mathsf{NP}$-$\mathsf{SRPT}$ is that, the average response time incurred under Poisson arrival is asymptotic optimal when the traffic intensity goes to $1$, if task service times are finite or job size distribution satisfies some mild conditions.
@@ Acknowledgments
We thank Ziv Scully for his many invaluable suggestions, including simplifying the proof for bounding the remaining workload under $\mathsf{NP}$-$\mathsf{SRPT}$, and the usage of the WINE identity. We also appreciate his feedback on improving the presentation of earlier versions of this paper.
Appendix A Proof of Lemma 10
Proof: In the following of the proof, we use $\mathcal{C}^{[k]}(\pi,t)=\cup_{i=1}^{k}\mathcal{C}_{i}(\pi,t)$ to denote the collection of jobs in the first $k$ classes, and let $W_{\pi}^{[k]}(t)=\sum_{i=1}^{k}{W_{\pi}^{(i)}(t)}$ represent the total remaining workload of jobs in the first $k$ classes, where $W_{\pi}^{(k)}(\pi,t)$ denotes the amount of remaining workload of jobs in class $\mathcal{C}_{k}(\pi,t)$. $W_{\pi^{*}_{1,N}}^{(k)}(t)$ and $W_{\pi^{*}_{1,N}}^{[k]}(t)$ are defined in a similar way for $\pi^{*}_{1,N}$. Without loss of generality, we assume $\log p_{\max}$ and $\log p_{\min}$ are integers.
For $\forall t\geq 0$, the number of unfinished jobs under the optimal algorithm is no less than,
$n_{\pi^{*}_{1,N}}(t)\geq$ $\sum_{k=\log p_{\min}}^{\log p_{\max}+1}{\frac{W_{\pi^{*}_{1,N}}^{(k)}(t)}{2^{k}}}$
$=$ $\sum_{k=\log p_{\min}}^{\log p_{\max}+1}{\frac{\Big{[}W_{\pi^{*}_{1,N}}^{[k]}(t)-W_{\pi^{*}_{1,N}}^{[k-1]}(t)\Big{]}}{2^{k}}}$
$=$ $\frac{W_{\mathrm{\pi^{*}_{1,N}}}^{[\log p_{\max}+1]}(t)}{2^{\log p_{\max}+1}}+\sum_{k=\log p_{\min}}^{\log p_{\max}+1}{\frac{W_{\mathrm{\pi^{*}_{1,N}}}^{[k]}(t)}{2^{k+1}}}$
$\geq$ $\sum_{k=\log p_{\min}}^{\log p_{\max}+1}{\frac{W_{\pi^{*}_{1,N}}^{[k]}(t)}{2^{k+1}}}.$ (16)
On the other hand, the number of jobs alive under $\mathsf{NP}$-$\mathsf{SRPT}$ can be upper bounded in a similar fashion,
$n_{\mathsf{NP-SRPT}}(t)\leq$ $\sum_{k=\log p_{\min}}^{\log p_{\max}+1}{\frac{W_{\mathsf{NP-SRPT}}^{(k)}(t)}{2^{k-1}}}$
$=$ $\sum_{k=\log p_{\min}}^{\log p_{\max}+1}{\frac{\Big{[}W_{\mathsf{NP-SRPT}}^{[k]}(t)-W_{\mathsf{NP-SRPT}}^{[k-1]}(t)\Big{]}}{2^{k-1}}}$
$=$ $\sum_{k=\log p_{\min}}^{\log p_{\max}}{\frac{W_{\mathsf{NP-SRPT}}^{[k]}(t)}{2^{k}}}+\frac{W_{\mathsf{NP-SRPT}}^{[\log p_{\max}+1]}(t)}{2^{\log p_{\max}}}$
$\leq$ $\sum_{k=\log p_{\min}}^{\log p_{\max}+1}{\frac{W_{\mathsf{NP-SRPT}}^{[k]}(t)}{2^{k-1}}}.$
Using Lemma 9, we are able to relate the number of unfinished jobs under two algorithms,
$n_{\mathsf{NP-SRPT}}(t)\leq$ $\sum_{k=\log p_{\min}}^{\log p_{\max}+1}{\frac{W_{\mathsf{NP-SRPT}}^{[k]}(t)}{2^{k-1}}}$
$\leq$ $\sum_{k=\log p_{\min}}^{\log p_{\max}+1}{\frac{W_{\pi^{*}_{1,N}}^{[k]}(t)}{2^{k-1}}}+\sum_{k=\log p_{\min}}^{\log p_{\max}+1}{\frac{N\cdot(2^{k}+\eta)}{2^{k-1}}}$
$\leq$ $4n_{\pi^{*}_{1,N}}(t)+4N\cdot\Big{(}\log\alpha+\beta+1\Big{)},$
where the last inequality follows from inequality (16). $\Box$

### v2

@@ VII Experimental Results
VIII Conclusion
In this work, we study the multitask scheduling problem, for which the optimal algorithms and tight analyses remain widely open for almost all settings. We propose $\mathsf{NP}$-$\mathsf{SRPT}$ algorithm, which achieves a competitive ratio that is order optimal when the number of machines is constant. Another appealing and more important property of $\mathsf{NP}$-$\mathsf{SRPT}$ is that, the average response time incurred under Poisson arrival is asymptotic optimal when the traffic intensity goes to $1$, if task service times are finite or job size distribution satisfies some mild conditions. Furthermore, by extending the analysis to scenarios with unknown job sizes, we demonstrate the robustness of our approach and establish the asymptotic optimality of policies for such settings.
@@ Acknowledgments
The author thanks Ziv Scully for his many invaluable suggestions, including simplifying the proof for bounding the remaining workload under NP-SRPT and the usage of the WINE identity, and for his feedback on improving the presentation of earlier versions of this article. The author also thanks the anonymous reviewers for their valuable comments and suggestions.